OpenStudy (calculusxy):

MEDAL! I need help on transformation \(y = x^2 - 8x + 12\)

OpenStudy (calculusxy):

@satellite73

OpenStudy (calculusxy):

@SolomonZelman

OpenStudy (anonymous):

transform in to what?

OpenStudy (anonymous):

vertex form maybe?

OpenStudy (calculusxy):

no it's how it will transform from y = x^2

OpenStudy (anonymous):

same thing as writing in vertex form

OpenStudy (calculusxy):

my vertex form was \((x - 4)^2 -4 \)

OpenStudy (anonymous):

half of 8 is 4 so it it going to look like \[y=(x-4)^2+k\] to find \(k\) replace \(x\) by \(4\)

OpenStudy (anonymous):

ok i believe you

OpenStudy (anonymous):

that means it is shifted how?

OpenStudy (calculusxy):

i know that the a part is whether the parabola faces up or down (is there is a mathematical term for this?)

OpenStudy (anonymous):

you have two \(-4\)'s there the one inside means shifted RIGHT 4, the one outsides means DOWN 4

OpenStudy (anonymous):

it faces up because the leading coefficient is positive (it is 1)

OpenStudy (calculusxy):

i know that the a part is whether the parabola faces up or down (is there is a mathematical term for this?)

OpenStudy (calculusxy):

the h part (horizontal translation) do i have to look at -4 or 4 (the one that will replace x to get 4 - 4 = 0 ?

OpenStudy (anonymous):

the vertex of \(y=x^2\) is \((0,0)\) and the vertex of \(y=(x-4)^2-4\) is \((4,-4)\) so you see the shift a) right 4 b) down 4

OpenStudy (anonymous):

i guess you should think "look at the 4" in the\((x-4)^2\) part right 4, not left 4

OpenStudy (anonymous):

so yes, the one that you replace x by to get 0

OpenStudy (calculusxy):

so i look at -4?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

vertex is \((4,-4)\)

OpenStudy (calculusxy):

okay. so that means that the horizontal translation will move 4 spaces to the left and the vertical translation will move 4 spaces down (since k is -4)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

horizontal translation os 4 units to the RIGHT

OpenStudy (anonymous):

how do you get from the vertex of \(y=x^2\) with is \((0,0)\) to the vertex of \(y=(x-4)^2-4\) which is \((4,-4)\)?

OpenStudy (calculusxy):

can u give me the rules for horizontal translation?

OpenStudy (anonymous):

i think of it like this compared to the graph of \(y=f(x)\) the graph of \(y=f(x+h) \) is shifted LEFT if \(h\) is positive, like \(f(x+3)\) would shift it 3 units left it is shifted RIGHT if \(h\) is negative, for example \(f(x-2)\) is shifted 2 units right

OpenStudy (calculusxy):

okay i got it. so the transformations are: 1) the parabola is going to face up 2) the horizontal translation is 4 spaces to the right because it is positive 4 3) the vertical translation is 4 spaces down because it is -4

OpenStudy (anonymous):

for quadratics some books write it like this: the vertex of \[y=a(x-h)^2+k\] is \((h,k)\)

OpenStudy (anonymous):

right, what you said

OpenStudy (calculusxy):

thanks!

OpenStudy (anonymous):

yw